# SADA: Safe and Adaptive Inference with Multiple Black-Box Predictions

### v1

@@ Abstract
Real-world applications often face scarce labeled data due to the high cost and time requirements of gold-standard experiments, whereas unlabeled data are typically abundant. With the growing adoption of machine learning techniques, it has become increasingly feasible to generate multiple predicted labels using a variety of models and algorithms, including deep learning, large language models, and generative AI. In this paper, we propose a novel approach that safely and adaptively aggregates multiple black-box predictions with unknown quality while preserving valid statistical inference. Our method provides two key guarantees: (i) it never performs worse than using the labeled data alone, regardless of the quality of the predictions; and (ii) if any one of the predictions (without knowing which one) perfectly fits the ground truth, the algorithm adaptively exploits this to achieve either a faster convergence rate or the semiparametric efficiency bound. We demonstrate the effectiveness of the proposed algorithm through experiments on both synthetic and benchmark datasets.
Keywords: Prediction-powered inference; multiple predictions; faster convergence rate; semiparametric efficiency bound
@@ 1 Introduction
In real-world applications, labeled data are often expensive, time-consuming, or requires expert knowledge to obtain, whereas unlabeled data are abundant and easier to collect. With the rapid advancement of machine learning and generative AI, generating predicted labels using powerful tools such as large language models (LLMs) and other algorithms has become easier than ever before. With predictions derived from a single source, many methods have been proposed, such as self-training (Lee,, 2013; Zhu et al.,, 2023) and prediction-powered inference (PPI) (Angelopoulos et al.,, 2023, 2024), to obtain better predictions or inference results under a variety of conditions.
However, the availability of multiple predictions introduces practical challenges. Outputs from different models–such as GPT, Llama, or DeepSeek–often differ, sometimes substantial; and the quality of predictions from black-box models can be highly variable. In particular, low-quality or poorly calibrated predictions can introduce significant noise, increasing variance and leading to unreliable inference. While one can evaluate the quality of each prediction or identify which model’s output is closest to the ground truth, a more compelling goal is to develop a principled method to aggregate multiple sources of predictions such that the resulting inference is guaranteed to perform no worse than using the labeled data alone. Moreover, if any one of the predictions is perfectly accurate or satisfies some ideal conditions, the aggregation method should achieve performance equivalent to using that prediction alone. Thus, this paper aims to answer the following question:
Given multiple predicted labels with unknown quality, how can we aggregate them in a safe and data-adaptive manner to improve estimation and inference?
To answer this question, we introduce a novel method that can safely and adaptively aggregate predictions from multiple black-box models, enabling valid and more informative inference. The proposed method has two key highlights:
  • Safety: The proposed method is guaranteed to perform no worse than the naive estimator (using the labeled data alone) in terms of mean squared error, regardless of the choice of machine learning models or their prediction accuracy. This means the method remains valid even when the ML predictions are arbitrarily misspecified.
  • Adaptivity: With multiple predictions, the proposed method adaptively utilizes more information of the better predictions to reduce variance, while downweighting poor predictions to avoid variance inflation. In particular, when one prediction is perfectly accurate, the method can effectively pick it up and performs as if we knew the ground truth of the unlabeled data.
1.1 Our novel contributions
First, we consider a semi-supervised setting where multiple sets of predicted labels are available, without making any assumptions about their quality or requiring prior knowledge of which predictions are more accurate. The predicted labels are also not needed to share the same scale or format, either with each other or with the true labels. This bridges the gap between advanced machine learning tools and principled methods for leveraging them to improve the inference results.
Second, we propose a safe and adaptive approach for aggregating multiple black-box predictions. Our method assigns data-driven weights to the predictions, effectively leveraging helpful information to reduce bias while mitigating the influence of harmful information that could increase variance. Importantly, it is guaranteed to perform no worse than using the labeled data alone, regardless of the quality of black-box predictions.
Third, we demonstrate that if any one of the predictions (no need to know which one) is perfectly accurate, the proposed estimator achieves a faster convergence rate, same as the oracle estimator that knows the ground truth of the unlabeled data. In addition, if we restrict the predictions to be deterministic functions of the observed features, say, generated by some pre-trained learners, and if any one of the predictions (no need to know which one) satisfies some ideal conditions, then the proposed estimator achives the best possible estimation efficiency among all regular asymptotically linear estimators; i.e., it attains the semiparametric efficiency bound.
@@ 1.2 Related work
Semi-supervised learning (SSL)
SSL has become a prominent approach in machine learning and statistics, leveraging both limited labeled data and abundant unlabeled data to enhance the performance of models (Grandvalet and Bengio,, 2004; Zhu,, 2005; Pan and Yang,, 2010).
@@
Over the past two decades, a wide range of SSL algorithms has been proposed. These methods differ in the assumptions, in how they utilize unlabeled data, and in their relationship to supervised learning approaches (van Engelen and Hoos,, 2020). Broadly, these algorithms can be categorized into two types: inductive and transductive methods. Inductive methods, similar to supervised learning, use a pre-trained model to assign labels to unlabeled data. Examples include self-training (Yarowsky,, 1995; Lee,, 2013; Berthelot et al.,, 2019, 2020), co-training (Blum and Mitchell,, 1998; Wang and Zhou,, 2010; Deng and Guo,, 2011), pseudo-labeled boosting methods (Zhou,, 2012), unsupervised preprocessing (Sheikhpour et al.,, 2017). In contrast, transductive methods do not produce a generalizable model; instead, they predict labels by directly propagating information through connections between data points. It typically defines a graph over all data points, both labeled and unlabeled, encoding the pairwise similarity of data points with possibly weighted edges (Jebara et al.,, 2009; Liu et al.,, 2012; Subramanya and Talukdar,, 2014). Additionally, significant progress has been made in recent years to understand, as well as how to leverage, the statistical benefits of the unlabeled data. Chakrabortty and Cai, (2018) and Azriel et al., (2022) studied linear regression problems within the SSL framework and proposed estimators that are more efficient than ordinary least squares (OLS) which relies solely on labeled data. Song et al., (2023) further extended this framework to general M-estimation problems. The methodology has also been adapted to high-dimensional settings, where the number of features exceeds the sample size (Zhang et al.,, 2019; Cai and Guo,, 2020; Zhang and Bradic,, 2022; Deng et al.,, 2024). Applications of SSL have expanded beyond statistical models to include both 2D computer vision tasks (Jeong et al.,, 2019; Liu et al.,, 2021; Tang et al.,, 2021; Zhou et al.,, 2022) and 3D object detection problems (Wang et al.,, 2021; Park et al.,, 2022; Li et al.,, 2023; Liu et al.,, 2023).
Prediction-powered inference (PPI)
@@
In the past few years, a growing body of research has focused on enhancing statistical inference by incorporating predictions from black-box AI/ML models (Wang et al.,, 2020; Motwani and Witten,, 2023). In particular, Angelopoulos et al., (2023) introduced prediction-powered inference (PPI), a framework that enables valid inference even when the predictive model is of low quality. However, PPI might perform worse in estimation efficiency compared to the naive method that uses labeled data only. This limitation has motivated further research aimed at improving the efficiency of PPI or integrating it with ideas from other statistical and machine learning frameworks. Examples include PPI++ (Angelopoulos et al.,, 2024), cross PPI (Zrnic and Candès,, 2024), stratified PPI (Fisch et al.,, 2024), and recalibrated PPI (Ji et al.,, 2025).
@@
In related work, Zhu et al., (2023) proposed a doubly robust self-training method that achieves faster convergence rates when predictions are highly accurate. Miao et al., (2024) introduced a post-prediction adaptive inference approach that ensures valid statistical inference without relying on assumptions about the ML predictions. Gan et al., (2024) explored a broader class of imputed loss functions to enhance modeling flexibility and efficiency. Gronsbell et al., (2025) focused on inference under squared error loss, situating PPI within the broader context of semiparametric theory. Bartolomeis et al., (2025) introduced a framework that integrates the predictions from multiple foundation models with randomized experiments while preserving valid statistical inference.
Missing data and causal inference
@@
SSL is also closely related to missing data and causal inference (Rubin,, 1974, 1976). In those problems, common approaches include likelihood-based inference (Dempster et al.,, 1977; Ibrahim,, 1990), imputation methods (Rubin and Schenker,, 1986; Rubin,, 2004; Vach and Schumacher,, 1993), Bayesian approaches (Rubin,, 1976), and semiparametric methods (Robins et al.,, 1994; Zhao et al.,, 1996). For more complex missing-not-at-random scenarios, earlier work established identification under specific modeling assumptions, such as outcome-selection models (Heckman,, 1979), pattern-mixture parametrizations (Little,, 1993, 1994), graphical models (Fay,, 1986; Ma et al.,, 2003), and sensitivity analysis techniques (Rotnitzky et al.,, 1998; Robins et al.,, 2000).
2 Problem Setup
@@
We first introduce some notations we use throughout. All vectors are assumed to be column vectors unless otherwise specified. Let $\mathbf{u}^{\otimes 2}=\mathbf{u}\mathbf{u}^{\mathrm{\scriptscriptstyle T}}$ for a vector $\mathbf{u}$. We denote $\mathbf{A}\preceq\mathbf{B}$ for two symmetric square matrices $\mathbf{A}$ and $\mathbf{B}$ if $\mathbf{B}-\mathbf{A}$ is positive semi-definite. We generally use capital letters to denote random variables, and the corresponding lowercase letters to denote their realizations. We use $\mathbb{P}$ to denote the probability measure and $\mathbb{E}$ to denote the expectation. For two random vectors $\mathbf{U}$ and $\mathbf{V}$, let $\operatorname{cov}(\mathbf{U},\mathbf{V})=\mathbb{E}[\{\mathbf{U}-\mathbb{E}(\mathbf{U})\}\{\mathbf{V}-\mathbb{E}(\mathbf{V})\}^{\mathrm{\scriptscriptstyle T}}]$ and $\operatorname{var}(\mathbf{U})=\operatorname{cov}(\mathbf{U},\mathbf{U})$.
SSL, objective and the naive estimator
We consider a standard SSL setting that consists of a set of $n$ labeled samples, $\mathcal{L}=\{(\mathbf{x}_{i},y_{i}),i=1,\dots,n\}$ and a set of NnN-n unlabeled samples, $\mathcal{U}=\{\mathbf{x}_{i},i=n+1,\dots,N\}$, drawn from some underlying distribution of random variables $(\mathbf{X},Y)$. Let $\ell_{\bm{\theta}}(\mathbf{x},y)$ be a convex loss function. In this paper, we focus on a $p$-dimensional parameter, ${\bm{\theta}}^{*}\in\Theta\subset\mathbb{R}^{p}$, defined as
${\bm{\theta}}^{*}=\operatornamewithlimits{arg\,min}_{{\bm{\theta}}\in\Theta}\mathbb{E}\{\ell_{\bm{\theta}}(\mathbf{X},Y)\}.$
We define the score function $\mathbf{s}(\mathbf{x},y;{\bm{\theta}})=\partial\ell_{\bm{\theta}}(\mathbf{x},y)/\partial{\bm{\theta}}$. Then, we can write $\bm{\theta}$$^{*}$ as the solution to the estimating equation
$\mathbb{E}\{\mathbf{s}(\mathbf{X},Y;{\bm{\theta}})\}=\mathbf{0}.$ (1)
The definition of $\bm{\theta}$$^{*}$ is intentionally broad and includes many commonly studied parameters. For example, if the goal is to estimate the outcome mean $\mathbb{E}(Y)$, one may define the loss function as $\ell_{\theta}(\mathbf{x},y)=(y-\theta)^{2}/2$, which yields the score function $\mathbf{s}(\mathbf{x},y;\theta)=y-\theta$. As another example, when the parameter of interest is the coefficient in a linear regression model, the loss can be defined as $\ell_{\bm{\theta}}(\mathbf{x},y)=(y-\mathbf{x}^{\mathrm{\scriptscriptstyle T}}{\bm{\theta}})^{2}/2$, leading to the score function $\mathbf{s}(\mathbf{x},y;{\bm{\theta}})=(y-\mathbf{x}^{\mathrm{\scriptscriptstyle T}}{\bm{\theta}})\mathbf{x}$.
The naive estimator, $\widehat{\bm{\theta}}^{\text{nv}}$, which only uses the labeled data, solves
Assume the Hessian matrix $\mathbf{H}=\mathbb{E}\{\partial\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}/\partial{\bm{\theta}}\}$ exists and is nonsingular. Under some regularity conditions (see Assumption 1), the mean squared error of $\widehat{\bm{\theta}}^{\text{nv}}$ equals
where ${\bm{\Sigma}}_{\text{nv}}=\operatorname{var}\{\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})\}$.
Availability of multiple predicted labels
We consider the availability of predicted labels from $K$ black-box models, denoted as $\widehat{\mathbf{y}}_{i}=(\widehat{y}_{1,i},\ldots,\widehat{y}_{K,i})$, for each subject $i=1,\ldots,N$. For example, in our application of wine reviews (see Section 5.2), a LLM can quickly predict a score based on information such as the wine’s name, price, vineyard, and other attributes. Multiple LLMs–such as GPT, Llama, and DeepSeek–can be used to generate a range of predictions. We make no assumptions on the quality of these predictions. The data structure and generating process are illustrated in Table 1 and left panel of Figure 1.

### v2

# SADA: Safe and Adaptive Aggregation of Multiple Black-Box Predictions in Semi-Supervised Learning

@@ Abstract
Semi-supervised learning (SSL) arises in practice when labeled data are scarce or expensive to obtain, while large quantities of unlabeled data are readily available. With the growing adoption of machine learning techniques, it has become increasingly feasible to generate multiple predicted labels using a variety of models and algorithms, including deep learning, large language models, and generative AI. In this paper, we propose a novel approach that safely and adaptively aggregates multiple black-box predictions of uncertain quality for both inference and prediction tasks. Our method provides two key guarantees: (i) it never performs worse than using the labeled data alone, regardless of the quality of the predictions; and (ii) if any one of the predictions (without knowing which one) perfectly fits the ground truth, the algorithm adaptively exploits this to achieve either a faster convergence rate or the semiparametric efficiency bound. We demonstrate the effectiveness of the proposed algorithm through small-scale simulations and two real-data analyses with distinct scientific goals. A user-friendly R package, sada, is provided to facilitate practical implementation.
Key Words: Black-box models; Prediction-powered inference; Semiparametric efficiency; Semi-supervised learning.
@@ 1 Introduction
In many scientific and industrial applications, acquiring gold-standard labels can be expensive, time-consuming, or require specialized expertise, while collecting unlabeled data is relatively cheap and scalable. This disparity motivates semi-supervised learning (SSL), a paradigm that combines scarce labeled observations with large amounts of unlabeled samples (Zhu, 2005; van Engelen & Hoos, 2020). Over the past decades, a wide range of SSL algorithms have been developed to exploit abundant unlabeled data to improve model performance, for example by using a pre-trained model to assign pseudo labels to unlabeled samples (Lee, 2013; Zhu et al., 2023) or by propagating information through relationships between data points (Jebara et al., 2009; Subramanya & Talukdar, 2014).
With the rapid evolution of machine learning and generative AI, it has become much easier than ever to generate predicted labels using powerful tools such as large language models (LLMs). These predictions can often be obtained at substantially lower cost and with much shorter turnaround time than human labels. Consequently, leveraging machine-learned predictions in SSL has become an increasingly important practical demand. In the context of statistical inference, Angelopoulos et al. (2023) introduced prediction-powered inference (PPI), a framework that enables valid inference without imposing any assumptions on the predictions. However, when the predictive model is of low quality, PPI may have worse estimation efficiency than labeled-only baselines. This limitation has motivated further work aimed at improving the efficiency or integrating it with ideas from other statistical and machine learning frameworks (Angelopoulos et al., 2024; Zrnic & Candès, 2024; Fisch et al., 2024; Ji et al., 2025; Miao et al., 2025).
Despite these advances, several gaps and challenges remain when one considers modern SSL pipelines in which multiple predicted labels are available. First, current PPI methods are designed for a single predictive model, whereas in practice it is common to obtain predictions from multiple models or sources. Although one may heuristically select the seemingly best model or apply ensembling strategies, there is no general framework within SSL for aggregating multiple predictions in a principled manner. Second, outputs from different models, such as GPT, Llama, or DeepSeek, can differ substantially, and the quality of predictions from black box models can be highly variable. In particular, low-quality or poorly calibrated predictions can introduce significant noise, increasing variance and leading to unreliable inference. Third, most existing work targets statistical inference for the paramter of interest; however, in many applications such as image classification, practitioners are primarily concerned with predictive model accuracy rather than parameter estimation. Current PPI methods and their extensions do not directly address how to improve prediction accuracy while retaining rigorous guarantees.
In this paper, we introduce SADA, a novel and comprehensive procedure that safely and adaptively aggregates multiple predicted labels originating from diverse black-box models for both inference and prediction tasks. Our approach is characterized by two key features: safety and adaptivity. Safety ensures that the proposed estimator never performs worse than a labeled-only baseline, regardless of the quality of the predictions. Specifically, for inference, the SADA estimator guarantees an asymptotic variance no larger than that of the estimator using labeled data only, a property that is generally not ensured by PPI (Angelopoulos et al., 2023) or PPI++ (Angelopoulos et al., 2024). For prediction, the excess risk of the SADA estimator converges at a rate no slower than that of the labeled-only baseline. Adaptivity means that our algorithm automatically aggregates multiple predictions by assigning them appropriate weights, leading to more efficient inference or improved predictive accuracy. Notably, in cases where one prediction is perfectly accurate, SADA automatically identifies and leverages it to attain a faster convergence rate or the semiparametric efficiency bound.
Our contributions are threefold. First, we propose a principled and fully data-driven procedure for aggregating multiple predicted labels of uncertain quality in a semi-supervised learning setting, accompanied by theoretical guarantees, which greatly enhances and broadens the applicability of PPI-based methods in practice. Second, we derive the asymptotic representations and excess risk bounds for the SADA estimator, which, make explicit how the method operates, demonstrate how it protects against potentially low-quality predictions, and clarify how it automatically borrows strength from informative predictions without requiring prior knowledge or formal testing of the quality of each model. Third, we illustrate these properties through intuitive explanations, simulation studies, and two real-data applications with distinct tasks that demonstrate the practical usefulness of the approach. We also provide a user-friendly R package, sada, to facilitate implementation of the proposed algorithm.
@@ 1.1 Related work
Semi-supervised learning (SSL).
@@
Over the past decades, a wide range of SSL algorithms has been proposed. These methods differ in the assumptions, in how they utilize unlabeled data, and in their relationship to supervised learning approaches (van Engelen & Hoos, 2020). Broadly, these algorithms can be categorized into two types: inductive and transductive methods. Inductive methods, similar to supervised learning, use a pre-trained model to assign labels to unlabeled data. Examples include self-training (Yarowsky, 1995; Lee, 2013; Berthelot et al., 2019, 2020), co-training (Blum & Mitchell, 1998; Wang & Zhou, 2010; Deng & Guo, 2011), pseudo-labeled boosting methods (Zhou, 2012), unsupervised preprocessing (Sheikhpour et al., 2017). In contrast, transductive methods do not produce a generalizable model; instead, they predict labels by directly propagating information through connections between data points. It typically defines a graph over all data points, both labeled and unlabeled, encoding the pairwise similarity of data points with possibly weighted edges (Jebara et al., 2009; Liu et al., 2012; Subramanya & Talukdar, 2014). Additionally, significant progress has been made in recent years to understand, as well as how to leverage, the statistical benefits of the unlabeled data. Chakrabortty & Cai (2018) and Azriel et al. (2022) studied linear regression problems within the SSL framework and proposed estimators that are more efficient than ordinary least squares (OLS) which relies solely on labeled data. Song et al. (2023) further extended this framework to general M-estimation problems. The methodology has also been adapted to high-dimensional settings, where the number of features exceeds the sample size (Zhang et al., 2019; Cai & Guo, 2020; Zhang & Bradic, 2022; Deng et al., 2024). Applications of SSL have expanded beyond statistical models to include both 2D computer vision tasks (Jeong et al., 2019; Liu et al., 2021; Zhou et al., 2022) and 3D object detection problems (Wang et al., 2021; Park et al., 2022; Liu et al., 2023).
Prediction-powered inference (PPI).
@@
In the past few years, a growing body of research has focused on enhancing statistical inference by incorporating predictions from black-box models (Wang et al., 2020; Motwani & Witten, 2023). In particular, Angelopoulos et al. (2023) introduced prediction-powered inference (PPI), a framework that enables valid inference even when the predictive model is of low quality. However, PPI might be less efficient than the labeled-only baseline, motivating follow-up works aimed at improving its efficiency or combining it with other statistical and machine learning ideas. Examples include PPI++ (Angelopoulos et al., 2024), cross PPI (Zrnic & Candès, 2024), stratified PPI (Fisch et al., 2024), recalibrated PPI (Ji et al., 2025).
@@
In related work, Zhu et al. (2023) proposed a doubly robust self-training method that achieves faster convergence rates when predictions are highly accurate. Miao et al. (2025) introduced a post-prediction adaptive inference approach that ensures valid statistical inference without relying on assumptions about the predictions. Gan et al. (2024) explored a broader class of imputed loss functions to enhance modeling flexibility and efficiency. Gronsbell et al. (2025) focused on inference under squared error loss, situating PPI within the broader context of semiparametric theory. Bartolomeis et al. (2025) introduced a framework that integrates the predictions from multiple foundation models with randomized experiments while preserving valid statistical inference.
Missing data and causal inference.
@@
SSL is also closely related to missing data and causal inference (Rubin, 1974, 1976). When missingness is at random, a rich variety of methods has been conducted to handle this, such as likelihood-based inference (Ibrahim, 1990), imputation (Rubin, 2004), and semiparametric methods (Robins et al., 1994; Zhao et al., 1996). For more complex missing-not-at-random settings, identification and estimation methods are also well developed, such as outcome selection models (Heckman, 1979), graphical models (Fay, 1986; Ma et al., 2003), sensitivity analysis techniques (Robins et al., 2000), and approaches based on instrumental variables or shadow variables (Wang et al., 2014; Zhao & Shao, 2015; Zhao & Ma, 2022; Miao et al., 2024).
1.2 Paper organization
The rest of this paper is organized as follows. Section 2 introduces the problem setup and data structure. Section 3 presents the SADA approach for inference and develops its theoretical properties, while Section 4 adapts the procedure to the prediction task. Section 5 constructs a small-scale simulation study to evaluate the finite-sample performance of the proposed methods. Section 6 applies the proposed methods to two real-data analyses with distinct tasks. Section 7 concludes the paper. All proofs are provided in the Supplementary Material.
1.3 Notations
@@
The following notations are used throughout the paper. All vectors are assumed to be column vectors unless otherwise specified. Let $\mathbf{u}^{\otimes 2}=\mathbf{u}\mathbf{u}^{\mathrm{\scriptscriptstyle T}}$ for a vector $\mathbf{u}$. We denote $\mathbf{A}\preceq\mathbf{B}$ for two symmetric square matrices $\mathbf{A}$ and $\mathbf{B}$ if $\mathbf{B}-\mathbf{A}$ is positive semi-definite. We generally use capital letters to denote random variables, calligraphic letters to denote their support, and lowercase letters to denote their realizations. We use $\mathbb{P}$ to denote the probability measure and $\mathbb{E}$ to denote the expectation. For two random vectors $\mathbf{U}$ and $\mathbf{V}$, let $\operatorname{cov}(\mathbf{U},\mathbf{V})=\mathbb{E}[\{\mathbf{U}-\mathbb{E}(\mathbf{U})\}\{\mathbf{V}-\mathbb{E}(\mathbf{V})\}^{\mathrm{\scriptscriptstyle T}}]$ and $\operatorname{var}(\mathbf{U})=\operatorname{cov}(\mathbf{U},\mathbf{U})$.
2 Setup
In this paper, we consider two types of scientific goals: inference and prediction. Let $Y$ be a scaler outcome and $\mathbf{X}$ be a vector of covariates. For the inference problem, the target is to estimate and draw inference on a $p$-dimensional parameter, ${\bm{\theta}}^{*}\in\Theta\subset\mathbb{R}^{p}$, defined through an estimating equation
$\mathbb{E}\{\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})\}=\mathbf{0},$ (1)
where $\mathbf{s}(\mathbf{x},y;{\bm{\theta}})$ is a user-specified $p\times 1$ vector-valued function. This formulation is broad and include many commonly studied parameters, such as outcome means, quantiles, least squares coefficients, and other functionals of $\mathbf{X}$ and $Y$, including the unique minimizer of a loss function or the maximizer of a criterion function (van der Vaart, 1998). The goal is to construct an estimator of $\bm{\theta}$$^{*}$ that is as close as possible to the true value $\bm{\theta}$$^{*}$.
By contrast, a prediction problem is defined by a hypothesis family $\{f_{\bm{\theta}}:\mathcal{X}\to\mathcal{Y}\}$ parameterized by ${\bm{\theta}}\in\Theta$, together with a loss function $\ell(f_{\bm{\theta}}(\mathbf{x}),y)$ that measures the discrepancy between the model’s output and the true label. For example, in a binary classification problem, the cross-entropy loss is given by $\ell(f_{\bm{\theta}}(\mathbf{x}),y)=-y\log\{f_{\bm{\theta}}(\mathbf{x})\}-(1-y)\log\{1-f_{\bm{\theta}}(\mathbf{x})\}$ with $f_{\bm{\theta}}(\mathbf{x})\in(0,1)$. Define $\mathcal{R}({\bm{\theta}})=\mathbb{E}\{\ell(f_{\bm{\theta}}(\mathbf{X}),Y)\}$ as the expected prediction loss, and
${\bm{\theta}}^{*}=\operatornamewithlimits{arg\,min}_{{\bm{\theta}}\in\Theta}\mathcal{R}({\bm{\theta}}),$
as the optimal prediction parameter. The goal is to construct an estimator $\widehat{\bm{\theta}}$ such that the excess risk, $\mathcal{R}(\widehat{\bm{\theta}})-\mathcal{R}({\bm{\theta}}^{*})$, is as small as possible.
We consider a standard semi-supervised learning setting in which only a subset of the sample is labeled with the ground-truth outcome $Y$. Denote the data as $\mathcal{L}\cup\mathcal{U}$, where $\mathcal{L}=\{(\mathbf{x}_{i},y_{i}),i=1,\dots,n\}$ is labeled set and $\mathcal{U}=\{\mathbf{x}_{i},i=n+1,\dots,N\}$ is the unlabeled set. Assume that $\mathcal{L}$ and $\mathcal{U}$ are independent and share the same marginal distribution of $\mathbf{X}$.
In addition, we consider the availability of predicted labels from $K$ black-box models, denoted by $\widehat{\mathbf{Y}}=(\widehat{Y}_{1},\ldots,\widehat{Y}_{K})^{\mathrm{\scriptscriptstyle T}}$. These predictions may originate from any low-cost source: crowdsourced annotators, weak-supervision pipelines, measurement-error surrogates, model ensembles, traditional machine learning models, or large pre-trained models whose quality may be uncertain. We do not require the prediction $\widehat{Y}_{k}$ to share the same magnitude or functional form, either with one another or with the true outcome $Y$. For example, when $Y$ is a binary label, each $\widehat{Y}_{k}$ may be either categorical or continuous. We allow the generation process of $\widehat{Y}_{k}$ to be a black box, potentially depending not only on the observed feature $\mathbf{X}$ but also on some latent or unobserved variables. Predicted labels are available for all subjects, both labeled and unlabeled, and are denoted by $\widehat{\mathbf{y}}_{i}=(\widehat{y}_{1,i},\ldots,\widehat{y}_{K,i})^{\mathrm{\scriptscriptstyle T}}$, for each subject $i=1,\ldots,N$. We assume that the distribution of $\widehat{\mathbf{Y}}$ is the same for the labeled and unlabeled samples. The data structure and overall generating process are illustrated in Table 1 and Figure 1.